Suppose X is a Binomial random variable for which there are 4 independent trials and probability of success 0.4. What is P(X > 0)?

a.

0.528

b.

0.1640

c.

0.8704

d.

0.4

e.

0.7638
Answer #1

n = 4

P = 0.4

It is a binomial distribution.

P(X = x) = nCx * px * (1 - p)n - x

P(X > 0) = 1 - P(X < 0)

= 1 - (4C0 * (0.4)^0 * (0.6)^4)

= 1 - 0.1296

= 0.8704

Option - c is correct.